OpenStudy (anonymous):

One number is 5 more than twice another and their sum is 23. Which of the following system of equations represents the word problem? A. x + y = 23 and y = 2x + 5 B. y = x + 5 and y + 2(x + 5) = 23 C. y = 2(x + 5) and x + y = 23

OpenStudy (anonymous):

A. x + y = 23 and y = 2x + 5

OpenStudy (anonymous):

@Khadeejah : Would u explain :D

OpenStudy (anonymous):

Here is mine : Let the first number y ,and the second number x Since that One number is 5 more than twice another y-5=2x ,y=2x+5 Since their sum is 23 y+x=23

OpenStudy (jiteshmeghwal9):

let one no. be x. Then,sec. no.=x+5 =>(x+5)+(x)=23 =>2x+5=23 =>2x=23-5 =>2x=18 =>x=18/2 =>x=9